OpenStudy (anonymous):

What is the solution of |2x-1|=5?

OpenStudy (anonymous):

if we take the positive charge then we find 2x-1=5 =>2x=5+1 =>x=6/2 x=3 if we take the negative charge then we find -(2x-1)=5 =>-2x+1=5 =>-2x=5-1 -2x=4 x=-2

OpenStudy (anonymous):

-2,3 :D!

OpenStudy (anonymous):

On the same track as you met.

OpenStudy (anonymous):

=D